Two point size bodies of same mass are knotted to a horizontal string one at the end, other at the mid point of it. The string is rotated in horizontal plane with the other end as centre. If T is the tension in the string b/w centre of circles and first body then the tension in the string b/w the two bodies is

Grade:12

1 Answers

Abhishek Gupta
7 years ago
 
Let : 1-)tension in outer string is T’  ,  2-) Length of string as 2L  3-) mass of each body as m
And we know that length of string for outer mass will 2L
After drawing FBD for outer mass:-
T’ = m*square of angular velocity * 2L
T’ = m*w2*2L
T’ = 2mLw2                            ------1
Similarly for inner mass:-
T – T’ = m*w2*L
T = 3mLw2                            ------2
So, tension in outer string is T’ where 
T’ = 2T/3    (Answer)